### v3

# INTERVENOR[Uncaptioned image]: Prompt the Coding Ability of Large Language Models with the Interactive Chain of Repairing

Abstract
This paper proposes INTERactiVE chaiN Of Repairing (INTERVENOR), which mimics human code repairing behavior (iteratively judging, rethinking, and repairing) and prompts the coding ability of regard Large Language Models (LLMs). Specifically, INTERVENOR employs two LLM based agents, Code Learner and Code Teacher, to play different roles in code repairing and work interactively to repair the generated codes. The Code Learner is asked to generate and repair code according to the instructions from the Code Teacher. The Code Teacher rethinks the code errors according to the corresponding feedback from compilers and iteratively generates the chain-of-repairing (CoR) to guide the code repairing process for Code Learner. Our experiments show that INTERVENOR outperforms baseline models and achieves about 13% and 4.5% improvements over GPT-3.5 in code generation and code translation tasks, respectively. Our further analyses show that CoR can illuminate the bug reasons and solution plans via natural language. With the feedback of code compilers, INTERVENOR can accurately identify the syntax errors and assertion errors and provide precise instructions to repair codes. All data and codes are available at https://github.com/NEUIR/INTERVENOR.
INTERVENOR[Uncaptioned image]: Prompt the Coding Ability of Large Language Models with the Interactive Chain of Repairing
[…]
1 Introduction
Large Language Models (LLMs), such as ChatGPT OpenAI (2022), have demonstrated extremely strong emergency ability and benefited lots of tasks, especially for code generation OpenAI (2023); Roziere et al. (2023); Wang et al. (2023). It greatly stimulates the efficiency and productivity of coding and software development Qian et al. (2023a). Existing code based models focus on pretraining language models on code corpus Muennighoff et al. (2023); Luo et al. (2023); Li et al. (2023b); Zheng et al. (2023) and using Chain-of-Thought (CoT) to prompt the coding ability of LLMs Wei et al. (2023); Huang et al. (2023); Li et al. (2023a). Nevertheless, forcing LLMs to directly generate the totally correct coding results is challenging even for skilled programmers in real-world scenarios Chen et al. (2023).
Figure 1: The Illustration of INTERVENOR. There are two agents in INTERVENOR, the teacher and student, whom collaborate to fix the code. The error messages are utilized as a kind of INTERVENORRefer to caption to alleviate the Degeneration-of-Thought (DoT) problem.
Recently, researchers focus on self-improving code generation methods using Self-Repair methods Olausson et al. (2023); Chen et al. (2023), which employ LLMs to execute codes and repair codes by themselves for improving the quality of generated codes. Moreover, multi-agent collaborative coding approaches Qian et al. (2023a); Dong et al. (2023) have also proven their effectiveness in handling difficult code tasks by prompting LLMs to play different roles, such as developers and testers. However, lots of bugs are difficult to find due to the cognitive inertia McGuire (1960)–overlooking the buggy codes that may not conform to their pre-existing coding thinking of LLMs. Thus, these agent based code refine methods Dong et al. (2023); Qian et al. (2023a) are highly dependent on the self-evaluation ability of LLMs and may face the Degeneration-of-Thought (DoT) problem Liang et al. (2023); Shinn et al. (2023).
This paper proposes INTERactiVE chaiN Of Repairing (INTERVENOR) to intervene the DoT problem in code repairing by incorporating the feedback from code compilers to facility the bug detection. However, there is a significant gap between compiler errors and task instructions Wang et al. (2022b) for code repair. Thus, INTERVENOR proposes an interactive Chain-of-Repairing (CoR) mechanism to bridge the gap from compiler feedback and repair instruction. Follow Dong et al. (2023), we design two agents Code Learner and Code Teacher to repair codes. As shown in Figure 1, the Code Learner is responsible for both initial code generation and iterative code repair, while the Code Teacher iteratively generates the CoR to guide the code repairing process for Code Learner. Different from Self-Debug Chen et al. (2023), Code Teacher leverages the bug report from compilers and the program generated by Code Learner to rethink and generate insightful error explanations and comprehensive planning on how to address the errors. This interactive code repair process will continue until the code learner successfully fixes all the code errors or reaches the predetermined maximum number of repair attempts. Besides, we collect the buggy codes that are generated from GPT-3.5 and build the CodeError dataset to facilitate the code repairing related tasks.
Our experiments demonstrate the effectiveness of INTERVENOR by outperforming the state-of-the-art code generation and code translation models. Notably, INTERVENOR also achieves about 13% and 4.5% improvements over the GPT-3.5 model OpenAI (2022) in code generation and code translation tasks, showing its ability to improve the quality of generated codes via iterative code repairing. Our further analyses illustrate that INTERVENOR is effective in leveraging the bug messages from code compilers, recognizing the reasons for code errors, and also providing correction planning in natural language. Thanks to our CoR mechanism, INTERVENOR avoids thinking by LLMs themselves and can accurately diagnose the buggy codes and correct the assertion errors and name/syntax errors even in more difficult code generation scenarios. Our CoR mechanism enables LLMs to avoid designing complex prompts for generating codes and achieve the best performance via only three-turn code repairing. INTERVENOR shows the potential in leveraging the feedback from environments or rule systems to evolve LLMs Olausson et al. (2023).
2 Related Work
The code generation tasks Chen et al. (2021); Austin et al. (2021); Zheng et al. (2023) aim to generate correct and executable code based on the given natural language description, which has drawn lots of attention from researchers. LLMs such as ChatGPT OpenAI (2022) and GPT-4 OpenAI (2023) have shown strong effectiveness in generating code of high quality. To enhance the coding ability of LLMs, existing work focuses on code-specific pretraining and performs exceptionally well in code generation tasks Roziere et al. (2023); Li et al. (2023b); Luo et al. (2023); Wang et al. (2023).
Some models focus on utilizing some prompting techniques to enhance the coding capabilities of LLMs. CodeCoT Huang et al. (2023) is inspired by Chain-of-Thought Wei et al. (2023) and prompts the quality of generated codes using Code-CoT and the Self-exam CodeCoT methods. LLMs are asked to craft the code and design a set of test cases to generate and polish the codes. Structured Chain-of-Thought (SCoT) Li et al. (2023a) further considers the program structure, such as sequences, branches, and loops, and prompts LLMs to generate intermediate reasoning steps with program structures. Nevertheless, forcing LLMs to directly generate the correct program results is challenging in the code generation task Chen et al. (2023).
To better polish the generated codes, existing work focuses on Self-Refine Olausson et al. (2023) and Self-Repair Chen et al. (2023) methods. The Self-Refine models improve the quality of generated codes LLMs by decoding multiple samples and then selecting the most suitable one based on certain criteria. To build tailored criteria, one approach is to execute the generated codes and select the most optimal one based on the execution results Ni et al. (2023); Zhang et al. (2022); Shi et al. (2022); Li et al. (2022). Another approach is to rerank multiple code solutions to select the final code Shi et al. (2022); Zhang et al. (2022); Chen et al. (2022); Inala et al. (2022). However, these methods need to consume a lot of computing resources to generate code candidates, which is inefficient Zhang et al. (2023).
The other research direction is utilizing an iterative code repair approach to further improve the quality of generated codes Zhang et al. (2023); Welleck et al. (2022); Madaan et al. (2023); Shinn et al. (2023). Self-Debug Chen et al. (2023) utilizes the explanations generated by LLMs to fix self-generated code, whereas Self-Repair Olausson et al. (2023) employs human-provided feedback for improvement. Self-Edit Zhang et al. (2023) uses error messages to refine the generated code, but it requires training an additional fault-aware editor, which involves additional complex tasks such as model design and construction of training data. Instead of directly incorporating feedback for code repair, INTERVENOR designs an additional agent to rethink the reasons for coder errors and generate the chain-of-repairing to teach the other agent to repair codes via natural language.
Moreover, the work Dong et al. (2023); Qian et al. (2023a) also designs a multi-agent collaborative approach to simulate the software development process and improve the efficiency of code generation. Nevertheless, these methods are highly dependent on the self-evaluation ability of LLMs and may face the Degeneration-of-Thought (DoT) problem Liang et al. (2023); Shinn et al. (2023). Unlike them, INTERVENOR focuses on the bug-fixing process and proposes a simple but effective solution, which utilizes external tools Qian et al. (2023b), such as the Python interpreter, to execute the code Xu et al. (2023); Qian et al. (2023b) and use the accurate bug report to facility the agent collaboration during the interactive code repairing.
Figure 2: Illustration of Our Interactive Chain of Repairing Model (INTERVENOR Refer to caption).
3 Methodology
In this section, we introduce INTERVENOR, which conducts an interactive program repairing using LLM collaboration. We first describe the preliminary of code repairing (Sec. 3.1) and then introduce our interactive Chain-of-Repairing (CoR) mechanism (Sec. 3.2).
3.1 Preliminary of Code Repairing
The code repairing models mainly focus on self-repair Olausson et al. (2023); Chen et al. (2023). These models usually consist of three steps, including code generation, code execution, and code explanation. Self-repair aims to use LLM itself to conduct the self-debug and self-execution processes and then iteratively repair codes. Nevertheless, programmers usually fail to recognize the code errors because of cognitive inertia, making the code execution more difficult by LLM itself.
Different from these self-repair models, INTERVENOR follows previous work Zhang et al. (2023); Wang et al. (2022a) to incorporate the feedback from compilers to prompt the code generation ability of LLMs. Instead of directly feeding the code bug message to LLMs, we design the interactive chain of repairing mechanism, which builds two agents to rethink and repair the code errors. The compiler serves as an INTERVENOR to avoid the Degeneration-of-Thought (DoT) problem during the interactive repairing process.
3.2 Interactive Chain-of-Repairing (CoR)
As shown in Figure 2, given the coding task, INTERVENOR aims to mimic the human bug repairing behavior by iteratively acquiring feedback from compilers, then generating bug reasons and solving plans in natural language, and finally fixing the program. Specifically, INTERVENOR employs two LLM based agents to play different roles in code repairing (Sec. 3.2.1). Then it conducts an interactive code repairing process with the agents (Sec. 3.2.2).
3.2.1 Agent Building
INTERVENOR involves the integration of two agents, Code Learner and Code Teacher, who work interactively to repair the generated codes. The Code Learner generates codes and repairs codes according to the instruction from Code Teacher. The Code Teacher focuses more on rethinking and illuminating the code errors for students. More agent building details are shown in Appendix A.4.
Code Learner. The code learner conducts two coding behaviors, including first time code generation and interactive code repairing.
In the first time code generation process, the code learner aims to produce the initial code version according to the requirements of the given coding task. Then we change the agent role for code repairing and use the instruction “You are a student assistant with excellent code repair capabilities” to activate the code repairing ability of LLMs. Such an instruction aims to stimulate the ability of LLMs to conduct execution by themselves, rethink the code errors according to the code repair instruction from the code teacher, and fix the bugs.
Code Teacher. The code teacher plays a critical role in interactive code repairing. It aims to generate the code error explanation and bug fix planning for the code learner to help it repair codes.
Specifically, we use the prompt “You are an experienced and insightful programming instructor, and you need to identify the bugs in the given code based on the error messages.” to activate the bug tracing ability and code diagnosis ability. Subsequently, the teacher will offer detailed repair suggestions and guidance to help the student better comprehend and resolve the errors in the code.
3.2.2 Interactive Code Repairing Workflow
INTERVENOR conducts an interactive code repairing process and we will introduce the communication among agents and code compiler.
[…]
4 Experimental Methodology
In this section, we describe the datasets, evaluation metrics, baselines, and implementation details in our experiments.
Benchmark Language Problems #Tests
HumanEval Python 164 7.8
MBPP Python 500 3.1
HumanEval-X C++ 164 7.8
Java 164 7.8
JavaScript 164 7.8
CodeError Python 4463 9.0
Table 1: Statistics of the Data Used in Our Experiments. #Tests represents the average number of test cases.
Dataset. We evaluate the code generation and translation effectiveness on three datasets, including HumanEval, MBPP and HumanEval-X. Besides, we build a new benchmark CodeError to further test the code repairing ability. All data statistics are shown in Table 1.
HumanEval Chen et al. (2021) serves as a benchmark for evaluating the functional correctness of synthesized programs generated from docstrings. It comprises 164 hand-written Python programming problems, which consist of function signatures, docstrings, bodies, and multiple unit tests. MBPP Austin et al. (2021) is a benchmark that includes 974 introductory-level Python programming problems. Each problem comprises a problem statement, a code solution, and three automated test cases, and the task IDs that range from 11 to 510 are used for evaluation. HumanEval-X  Zheng et al. (2023) is a benchmark that is used to assess a model’s multi-programming language generation and translation capability. It consists of 820 human-crafted data instances, covering C++, Java, JavaScript, and Go.
Then we build the CodeError benchmark to further evaluate the code repairing effectiveness of INTERVENOR. We collect data from HumanEval and MBPP and reserve the instances, which conduct code errors in the code generation experiment. Specifically, for these code generation tasks, we utilize GPT-3.5 to generate codes according to the task instructions. These generated codes usually contain various bugs and the code errors include syntax errors, logical errors, and others. In this paper, we use the Python interpreter to test the generated code and reserve the buggy codes that do not pass the Python interpreter. These buggy codes and error messages are used to construct the CodeError benchmark. We provide more details about the CodeError benchmark in Appendix  A.2.
Evaluation Metrics. We use $Pass@k$ Chen et al. (2021) to evaluate the code generation and translation effectiveness of different models, which is the same as the previous work Chen et al. (2021); Zheng et al. (2023); Li et al. (2023a); Chen et al. (2022); Nijkamp et al. (2023).
Baselines. We first compare INTERVENOR with several code-oriented large language pretrained models, such as Incoder Fried et al. (2022), CodeGen Nijkamp et al. (2023), CodeGeeX Zheng et al. (2023), CodeT5 Wang et al. (2023), StarCoder Li et al. (2023b), WizardCoder Luo et al. (2023), and Llama based models Touvron et al. (2023); Roziere et al. (2023). These models are pretrained on large-scale code corpora, demonstrating strong code generation capabilities. Additionally, we also compare INTERVENOR with some closed-source and high-performance large language models, e.g. Claude Anthropic (2023), GPT-3.5 OpenAI (2022), and GPT-4 OpenAI (2023), which show strong emergent abilities, especially for the code generation tasks. In our experiments, GPT-3.5 is our main baseline model. Besides, we also compare Self-Debug Chen et al. (2023) and the multi-agent collaborative method, Self-Collaboration Dong et al. (2023) to show the code repairing ability of INTERVENOR.
Implementation Details. In our experiments, we use GPT-3.5 as the foundation model to build different agents in our INTERVENOR model. We set the temperature to 0.2 and the maximum generation length to 512 tokens. The maximum number of interactive code repairs is set to 5.
5 Evaluation Results
In this section, we evaluate the overall performance of INTERVENOR. Then we conduct ablation studies and also show the effectiveness of Interactive CoR in different testing scenarios. Finally, case studies are presented.
Model Code Generation Code Translation
HumanEval MBPP HumanEval-X Target Language
Python Python C++ Java JS Python C++ Java JS
InCoder Fried et al. (2022) 15.2 19.4 9.5 9.1 13.0 - - - -
CodeGen Nijkamp et al. (2023) 18.3 20.9 18.1 14.9 18.4 40.7 37.6 35.4 51.8
CodeGeeX Zheng et al. (2023) 22.9 24.4 17.1 20.0 17.6 68.5 43.6 56.8 45.2
CodeT5+ Wang et al. (2023) 30.9 - - - - - - - -
InstructCodeT5+ Wang et al. (2023) 35.0 - - - - - - - -
PaLM-Coder Chowdhery et al. (2022) 35.9 47.0 - - - - - - -
StarCoder Li et al. (2023b) 40.8 49.5 - - - - - - -
WizardCoder Luo et al. (2023) 57.3 51.8 - - - - - - -
LLama2 Touvron et al. (2023) 30.5 45.4 - - - - - - -
CodeLLama Roziere et al. (2023) 62.2 61.2 - - - - - - -
PanGu-Coder2 Shen et al. (2023) 61.6 - - - - - - - -
Claude Anthropic (2023) 47.6 - - - - - - - -
GPT-4 OpenAI (2023) 67.0 - - - - - - - -
Self-Debug Chen et al. (2023) 73.8 - - - - - - - -
Self-Collaboration Dong et al. (2023) 74.4 68.2 - - - - - - -
GPT-3.5 OpenAI (2022) 60.3 39.8 52.4 50.6 54.3 84.3 71.5 81.7 84.6
INTERVENOR 75.6 69.8 67.1 68.3 67.1 89.8 75.6 85.4 88.3
Table 2: Overall Performance of Different Models. We evaluate model effectiveness on code generation and code translation tasks using the $Pass@1$ evaluation metric. The baseline results are borrowed from corresponding papers. More evaluation results on the code translation task are shown in Appendix A.1.
5.1 Overall Performance
The overall performance of INTERVENOR in code generation and translation tasks is shown in Table 2.
Overall, INTERVENOR outperforms all baselines in all tasks by achieving more than 1% improvements, showing its effectiveness. Compared to our main baseline model GPT-3.5, INTERVENOR achieved about 13% and 4.5% improvements in code generation and code translation tasks, respectively. It illustrates that INTERVENOR has the ability to prompt the coding ability of LLMs by mimicking the human code repairing behavior–iteratively judging, rethinking, and repairing. Notably, INTERVENOR also surpasses Self-Debug and Self-Collaboration models, demonstrating its ability to successfully intervene in the code generation/translation process and guide LLMs to better repair the codes using our chain of repairing mechanism. All these experimental results highlight the generalization ability of INTERVENOR in improving LLMs’ coding ability in different languages.
Code Repair Prompt Methods HumanEval MBPP HumanEval-X CodeError Avg.
Python Python C++ Java JS Python
No Repair Zero-Shot 60.3 39.8 52.4 50.6 54.3 - -
Zero-Shot CoT 51.8 35.2 48.2 45.7 45.4 - -
Few-Shot 62.2 45.4 53.1 62.2 43.3 - -
Few-Shot CoT 60.4 45.4 48.2 63.4 57.9 - -
Single Turn Zero-Shot 62.2 41.6 54.3 65.2 60.4 4.9 48.1
Few-Shot 65.2 40.6 57.9 65.2 62.8 9.8 50.3
CoT 66.5 48.8 56.7 62.2 60.4 10.3 50.8
Self-Refine 65.2 48.8 57.3 64.0 60.4 5.2 50.1
Error Msgs 67.1 51.8 57.3 59.8 62.8 9.8 51.4
INTERVENOR (CoR) 69.5 51.0 59.8 65.2 60.4 15.9 53.6
Multi-Turns INTERVENOR (CoR) 75.6 69.8 67.1 68.3 67.1 21.7 61.6
Table 3: Ablation Studies. We compare different prompting techniques to evaluate the effectiveness of our chain-of-repairing mechanism. All models are built based on GPT-3.5 and evaluated with $Pass@1$.
Code Learner Code Teacher HumanEval MBPP HumanExal-X CodeError
Python Python C++ Java JS Python
CodeLlama-7B N/A 32.3 34.4 30.3 29.8 33.5 -
CodeLlama-7B 32.9 35.1 31.0 31.0 34.1 4.6
CodeLlama-13B 33.5 35.3 32.2 31.5 35.5 4.9
GPT-3.5 36.6 38.4 33.5 32.3 40.4 8.6
CodeLlama-13B N/A 39.6 36.4 36.6 33.5 39.6 -
CodeLlama-7B 40.0 37.1 37.1 34.5 40.4 8.8
CodeLlama-13B 42.1 37.7 38.6 36.6 42.8 9.4
GPT-3.5 43.6 40.4 42.2 37.5 46.1 12.6
GPT-3.5 N/A 60.3 39.8 52.4 50.6 54.3 -
CodeLlama-7B 65.2 42.8 55.4 55.1 56.1 11.8
CodeLlama-13B 66.5 46.1 56.1 58.9 58.9 13.6
GPT-3.5 69.5 51.0 59.8 65.2 60.4 15.9
Table 4: Model Performance of Different Code Learner and Code Teacher Model Pairings. We only conduct one turn code repair in experiments. “N/A” represents the initial output of Code Learner without any intervention from the Code Teacher for code repairing, it reflects the initial generation results of Code Learner.
5.2 Ablation Studies
The ablation studies are conducted to show the effectiveness of the interactive CoR mechanism.
Firstly, we evaluate the code generation ability of LLMs using different code generation/repairing prompting methods. As shown in Table 3, we compare CoT Kojima et al. (2023), Few-Shot Chen et al. (2021), and Few-Shot CoT Wei et al. (2023) models in experiments, which prompt LLMs to better generate codes. These models try to generate natural language as the chain of coding thought (CoT) or provide some instances to demonstrate the coding task (Few-Shot). Then we compare different methods to generate the code repair instruction, including Self-Refine Madaan et al. (2023), Error Msgs, and Chain-of-Repairing (CoR). Self-Refine asks the LLMs to rethink the errors by themselves, while Error Msgs and CoR incorporate the code error messages from compilers. Msgs directly uses error messages to guide the code repair process.
The experimental results show that code repairing is more effective than directly prompting LLMs to generate codes. Even though we conduct different methods to directly prompt LLMs to generate better codes, we achieve only 3.5% improvements. On the contrary, the code repairing methods improve the quality of generated codes by achieving more than 9.7% improvements with only single-turn repairing. Then we compare different code repairing methods. Both Error Msgs and CoR thrive on the feedback from code compilers and achieve more than 1.3% improvements than Self-Refine, demonstrating that compilers can provide valuable signals to help LLMs better recognize the code bugs. Notably, CoR achieves the best performance among all code repairing models, which illustrates its effectiveness in teaching LLMs for code repairing. The CoR mechanism uses the error messages from code compilers to prompt LLMs, aiming to rethink the reasons for making errors and generate the instructions for repairing.
Then, as shown in Table 4, we investigate the impact of using different LLMs to build Code Learner and Code Teacher. When we use the same LLM to build the Code Learner, the performance of INTERVENOR improves as the capabilities of the Code Teacher are enhanced, showing the effectiveness of our CoR mechanism. This indicates that stronger models can conduct a more in-depth analysis of erroneous code and provide more accurate and relevant suggestions for code repair. Similarly, by keeping the same LLM to build the Code Teacher, a stronger Code Learner also conducts better code repair results. This indicates that a more powerful Code Learner has a stronger instruct following ability to better comprehend the CoR provided by the Code Teacher for code repair.
Figure 3: The Impact of Different Code Repairing Turns. HumanEval, MBPP, and HumanEval-X (HEX) are used to evaluate our INTERVENOR model.
(a) AssertionError.
(b) Other Error Types.
Figure 4: Code Repairing Performance on the CodeError Dataset. We repair the error codes with one single turn. The codes are divided into two groups to evaluate the code repairing effectiveness, including Assertion Errors and Others (AttributeError, NameError, RecursionError, SyntacError and TypeError).
5.3 Effectiveness of INTERVENOR in Different Testing Scenarios
In this subsection, we delve deeper into exploring the effectiveness of INTERVENOR in two testing scenarios: 1) validating the impact of different code repairing turns, and 2) evaluating the code repairing effectiveness on different code error types.
Figure 5: Case Studies. We provide two cases that showcase the effectiveness of the Chain of Repairing generated by INTERVENOR when fixing AttributeError and AssertionError, respectively.
As shown in Figure 3, the code generation performance is significantly improved during the iteratively repairing. After three turns, the INTERVENOR achieves almost the best performance on HumanEval and HumanEval-X datasets, showing the efficiency of our interactive chain-of-repairing mechanism. For a more difficult dataset MBPP, INTERVENOR achieves 30% improvements and still has some room to achieve further improvements, demonstrating its advantages in dealing with more difficult and realistic coding tasks.
Then we show the code repairing effectiveness on different error types in Figure 4. We use different prompt methods to stimulate LLMs to repair code errors and show their effectiveness on different types of code errors. Overall, INTERVENOR doubles the number of corrected code examples of baseline models on both AssertionError and other error types, showing its effectiveness in repairing the code errors. On the one hand, INTERVENOR significantly outperforms other methods on the AssertionError task. It illustrates that INTERVENOR can provide more precise guidance and identify errors with the help of the bug report derived from the failed testing case. On the other hand, INTERVENOR also shows a strong ability to correct other code errors. Our chain-of-repairing method also thrives on the error messages of the compiler, breaks the cognitive inertia of LLMs, and identifies the specific code error for repairing. All these phenomena show that the quality of code execution feedback is critical in repairing codes.
5.4 Case Studies
Finally, we select two case studies in Figure 5 to demonstrate the effectiveness of INTERVENOR. We also compare INTERVENOR with Self-Refine Madaan et al. (2023), which prompts LLMs to execute codes, recognize bugs, and generate instructions by themselves.
Overall, the feedback from compilers indeed helps to improve the accuracy of code repairing by providing more valuable instructions. In the first case, Self-Refine fails to fix the AttributeError and adds the “all()” function in codes. On the contrary, INTERVENOR successfully fixes the code, showing its effectiveness. It accurately analyzes the reason for bugs “call the islower() method on an integer object, which is not possible” and also provides a solution by suggesting to “check if the key is a string before calling the islower() method”. In the second case, the instruction provided by Self-Refine thinks the code seems to be correct and does not offer definite solutions, showing the cognitive inertia in debug–It is hard to debug the code written by ourselves. INTERVENOR shows its effectiveness in directly generating the reason of bug “the function is not handling the case where the input list has only one element correctly”. More cases are shown in Appendix A.3.
6 Conclusion
This paper proposes INTERVENOR, which regards the feedback of compilers as the INTERVENOR to facilitate the agent collaboration of rethinking and repairing. Our experiments show that INTERVENOR outperforms previous code generation/repairing models by accurately diagnosing the buggy codes with the CoR mechanism.
Limitations
The performance of INTERVENOR largely depends on the capabilities of the foundation model, requiring the model to have the ability to analyze code errors and repair code, which poses higher demands on the model’s capabilities. The gains brought by INTERVENOR with models of inferior capabilities might be limited. Moreover, for some codes with only logical errors, the compiler requires the assistance of test cases to provide feedback. The quality and coverage of test cases are crucial in this scenario. Inadequate or poorly designed test cases may not be sufficient to detect or help rectify these logical errors effectively. Therefore, the effectiveness of INTERVENOR also hinges on the robustness and comprehensiveness of the test cases used in conjunction with it.
Appendix A Appendix
A.1 Code Translation Results.
As shown in Table 5, we present more detailed evaluation results on the code translation task.
In general, INTERVENOR shows the best performance across all twelve cross-language code translation tasks, demonstrating its ability to understand and translate codes. Compared with GPT-3.5, INTERVENOR improves its performance on all code translation tasks, especially on the tasks that translate other languages into Python. This indicates that INTERVENOR is more effective in generating an effective code repair chain according to the bug reports from the Python interpreter. Such a phenomenon aligns with human intuition during the debugging process, namely that bugs in Python code are more easily resolved.
A.2 More Details of the CodeError Benchmark
In our experiments, we build the CodeError benchmark to evaluate the code repair capabilities of LLMs, which facilitates the research on the code repairing task. The CodeError dataset is collected from basic programming problems, data analysis problems, and programming competition problems. In this section, we show more detailed data information about the CodeError benchmark.
Figure 6: An Example of the CodeError Dataset. The CodeError benchmark asks LLMs to fix buggy codes and then evaluate whether the fixed code meets the requirements specified in the task description and passes all test cases. During the process of code repairing, LLMs can utilize error messages to repair codes.
Source Model Target
Python C++ Java JS
Python CodeGen - 35.9 29.3 43.4
CodeGeeX - 34.2 42.0 34.8
GPT-3.5 - 62.8 70.7 82.3
INTERVENOR - 67.7 75.1 87.8
C++ CodeGen 33.8 - 43.2 54.5
CodeGeeX 62.8 - 71.7 50.8
GPT-3.5 81.1 - 89.6 82.3
INTERVENOR 90.2 - 92.1 86.1
Java CodeGen 52.7 41.4 - 57.7
CodeGeeX 75.0 49.7 - 50.0
GPT-3.5 89.6 75.6 - 89.1
INTERVENOR 92.1 79.3 - 90.9
JS CodeGen 35.5 35.4 33.8 -
CodeGeeX 67.7 46.9 56.6 -
GPT-3.5 82.3 76.2 84.8 -
INTERVENOR 87.2 79.9 89.1 -
Table 5: Code Translation Performance. We evaluate the code translation effectiveness among different program languages, including Python, C++, Java, and JavaScript (JS). We report the results of INTERVENOR, which only repair code with a single turn. All evaluation results are evaluated with $Pass@1$.
Dataset Language Size Avg.TC Error Type Error Msg Category
DeepFix Yasunaga and Liang (2021) C 6,971 CE Only Basic
Review4Repair Huq et al. (2022) Java 2,961 All Basic
Bug2Fix Lu et al. (2021) Java 5,835 All Basic
Github-Python Yasunaga and Liang (2021) Python 15k CE Only Basic
FixEval Haque et al. (2023) Java/Python 43k/243k 25.0 All Competition
CodeError (Ours) Python 4,463 9.0 All
Basic
Data Analysis
Competition
Table 6: A Comparison between CodeError and Other Code Repair Benchmarks. Size only represents the size of the test set. Avg.TC indicates the average number of test cases per problem. CE indicates compilation errors (e.g., SyntaxError). Error Msg indicates whether the buggy code contains detailed error information, such as the line of incorrect code, the reasons for the error, etc. Category represents the scope covered by the buggy code, CodeError covers basic programming problems, data analysis problems, and programming competition problems.
Basic Comp DA Total/Avg.
Problem 326 3888 249 4463
AssertionError 236 2949 52 3237
NameError 22 62 32 116
TypeError 39 91 53 183
IndexError 1 92 8 101
ValueError 2 229 44 275
SyntaxError 11 375 4 390
Other Errors 15 90 57 162
Avg.Problem Words 10 47 140 49
Avg.Buggy Code 21 34 2 31
Avg.Test Cases 4 10 1.6 9
Table 7: Data Statistics of CodeError. We calculate the average word count per problem, the average number of lines in buggy code, and the average number of test cases per problem. Basic, Comp, and DA represent basic programming problems, programming competition problems, and data analysis problems, respectively.
AssertionError is an exception that is raised when an assert statement fails. In this paper, an AssertionError indicates that the code can run correctly but fails to pass certain test cases, suggesting the presence of potential logic errors that require further investigation and resolution.
NameError is an exception that is raised when a local or global name is not found. This error occurs when you try to access a variable or a function that is not defined or is not in the current scope.
TypeError is an exception that is raised when an operation or function is applied to an object of an inappropriate type. This typically occurs when you try to perform an operation that is not supported for the type of data you are working with.
IndexError is an exception that is raised when you try to access an index that does not exist in a list, tuple, or any other sequence. This typically happens when you attempt to access an index that is outside the range of the sequence.
ValueError is an exception that is raised when a built-in operation or function receives an argument that has the right type but an inappropriate value. Essentially, this error occurs when a function receives an argument of the correct type, but the value of the argument is not appropriate for the operation.
SyntaxError is an exception that is raised when there is an error in the syntax of your code. This can happen due to various reasons, such as missing parentheses, invalid keywords, or incorrect indentation.
Table 8: Common Code Errors and Their Descriptions.
CodeError Examples. As shown in Figure 6, each example in CodeError consists of a programming task description, a buggy code snippet, error messages, and test cases to evaluate the code correctness of the generated codes. If the fixed/generated code meets the requirements specified in the task description and passes all test cases, it proves that the repaired code is correct. During the process of code repairing, LLMs need to fix the buggy code and can also use the error messages from the buggy code for assistance. We provide more examples that are sampled from the basic programming problems, programming competition problems, and data analysis problems in Figure 7, Figure 8, and Figure 9, respectively.
Data Collection. CodeError is collected from different coding problems, making the dataset diverse and reliable.
To ensure the diversity of CodeError, we collect basic programming problems, data analysis problems, and programming competition problems from HumanEval, MBPP, DS-1000, APPS, and the programming contest site AtCoder. We conduct code generation experiments using GPT-3.5 and CodeLlama on the first four datasets, preserving the generated codes that contain errors. For the programming contest site AtCoder, we crawl real-world user-submitted buggy codes to ensure the reliability of CodeError. The diverse data sources allow us to build a comprehensive and robust dataset to estimate the code repairing ability of LLMs.
Data Statistics. The data statistics of CodeError are shown in Table  7.
The CodeError benchmark contains a total of 4,463 examples, evenly distributed across more than six different error types. These errors range from simple syntax errors to complex logic errors. And we provide detailed descriptions for common error types in Table 8. The predominant error type is the AssertionError (Logic Error). This phenomenon is quite normal in the real-world coding scenario since the current integrated development environment (IDE) can assist developers in avoiding simple errors such as SyntaxError and NameError but may not help to identify logic errors within the code. For each example, there are 9 test cases on average to evaluate the correctness of the repaired codes.
Dataset Comparison. Finally, we compare the CodeError dataset with other code repair benchmarks. The differences are shown in Table 6.
Similar to FixEval Haque et al. (2023), we utilize test cases to assess the functional correctness of the repaired code. Additionally, in comparison to other benchmarks Yasunaga and Liang (2021); Huq et al. (2022); Lu et al. (2021); Haque et al. (2023), we have two notable features. 1) We provide detailed error information for erroneous code, including error code localization, the cause of the error, error type, and more. 2) CodeError covers a variety of problem types, including basic programming problems, data analysis problems, and programming competition problems, enabling a more comprehensive benchmark to evaluate the code repairing capabilities of LLMs.
A.3 Additional Case Studies
In this subsection, we sample some cases from CodeError to demonstrate the effectiveness of the Chain of repairing (CoR) mechanism.
As shown in Figure 10, in the first case, there is a simple NameError in the buggy code, which indicates that ‘hashlib’ is not defined. We can see that the INTERVENOR recognizes hashlib as a Python package and provides a solution: “import the ‘hashlib’ module at the beginning of the code”. In the second case, the error in the code is more subtle, involving an operation between ‘int’ and ‘list’. The CoR explicitly states that this is a “valid operation” and provides a solution: “check if the current element in the list is a list or not before adding it to the sum”. With the help of the accurate instruction, INTERVENOR successfully repairs the code. In Figure 11, we demonstrate that INTERVENOR can effectively address the boundary issues, which are often quite tricky. INTERVENOR accurately identifies the code error in the loop condition and suggests a modification method: “modify the loop to iterate up to $n-1$ instead of $n$”. In Figure 12, we demonstrate that INTERVENOR can fix multiple error types in the code by iteratively repairing. Over two iterations, INTERVENOR identifies and accurately modifies the ValueError and TypeError in the code.
All these cases demonstrate the effectiveness of the chain of repairing (CoR) mechanism and highlight the effectiveness of CoR in identifying errors and providing appropriate solutions for code repairing.
Figure 7: An Example from the Basic Programming Problems.
Figure 8: An Example from the Programming Competition Problems. The error information includes the line of code where the error occurs and the test case that failed.
Figure 9: An Example from the Data Analysis Problem. We use the completion mode in DS-1000.
Figure 10: Case Studies. We present two cases from CodeError. The chain of repairing clearly identifies the NameError and TypeError in the buggy code and provides the correct modification methods, which are highlighted in red.
Figure 11: Handling Boundary Issues.
Figure 12: Handling Multiple Bugs.
Figure 13: Role Instructions in INTERVENOR. Within INTERVENOR, there are two LLM-based agents Code Teacher and Code Learner. We utilize specific instructions to ensure that they play the correct roles and carry out the intended tasks.
Figure 14: Zero-Shot, Zero-Shot CoT, and Few-Shot Prompts for Code Generation Tasks.
Figure 15: Few-Shot CoT Prompts for Code Generation Tasks.
Figure 16: Zero-Shot and Few-Shot Prompts for Code Repairing.
Figure 17: Few-Shot CoT Prompts for Code Repairing.
Figure 18: Prompts Used in Self-Refine Model. The instructions for repair plan generation and code repair are shown.
Figure 19: Prompts Used in Error Msgs Model. It conduct code Repairing directly use the code error messages.
A.4 Prompts for Agent Building
In this subsection, we show the prompts used in our experiments.
We first show the role instructions used by Code Learner and Code Teacher in Figure 13. We utilize specific instructions to ensure that the Code Teacher and Code Learner within INTERVENOR play the correct roles and carry out code inspection and code repair. Then, we provide prompts used in code generation and code repair. In Figures 14 to 15, we show the Zero-Shot, Zero-Shot CoT, Few-Shot, and Few-Shot CoT prompts for code generation tasks. In Figures 16 to 19, we show some code repair prompts, including Zero-Shot, Few-Shot, Few-shot CoT, Self-Refine, and Error Msgs. Self-Refine asks the LLMs to rethink the errors by themselves, while Error Msgs incorporate the code error messages from compilers and directly use them to guide the code repair process.